OpenStudy (anonymous):

The density of osmium (the densest metal) is 22.57 g/cm^3. If a 1.00-kg rectangular box of osmium has two dimensions of 4.00cm X 4.00cm, calculate the third dimension of the block.

OpenStudy (aaronq):

find the V=w*l*h since, d=m/V -> d=m/(w*l*h)

OpenStudy (anonymous):

is 4x4x4 the lengthxwidthxheight? or do I have to find the height honestly I don't even understand the question itself

OpenStudy (aaronq):

yeah, you have to find height

OpenStudy (anonymous):

thanks! I checked the back and got the right answer :)

OpenStudy (aaronq):

nice !